OpenStudy (anonymous):

would you set this equation to zero? x^2 +5x=24

OpenStudy (anonymous):

if you subtract 24 from both sides you are left with 0 on the right

OpenStudy (campbell_st):

if you want to solve it... subtract 24 from both sides then it can be factorised to find the solution.

OpenStudy (anonymous):

then you need to factor and solve or use the quadratic

OpenStudy (anonymous):

x^+5x-24=24-24 x^+5x-24=0 (x-8)(x+3)=0

OpenStudy (campbell_st):

correct, well done

OpenStudy (anonymous):

then I solve for x?

OpenStudy (anonymous):

x=8, x=-3

OpenStudy (campbell_st):

almost correct. you have a slight error in the factorisation. the larger factor is positive since the middle term is + 5x.. so reverse the signs in the factors

OpenStudy (anonymous):

gotcha! x=-8, x=3

OpenStudy (campbell_st):

thats better and thats the solution , well done

OpenStudy (anonymous):

you are awesome! thank you :)